# Accelerating Targeted Hard-Label Adversarial Attacks in Low-Query Black-Box Settings

### v1

@@ Abstract
Deep neural networks for image classification remain vulnerable to adversarial examples – small, imperceptible perturbations that induce misclassifications. In black-box settings, where only the final prediction is accessible, crafting targeted attacks that aim to misclassify into a specific target class is particularly challenging due to narrow decision regions. Current state-of-the-art methods often exploit the geometric properties of the decision boundary separating a source image and a target image rather than incorporating information from the images themselves. In contrast, we propose Targeted Edge-informed Attack (TEA), a novel attack that utilizes edge information from the target image to carefully perturb it, thereby producing an adversarial image that is closer to the source image while still achieving the desired target classification. Our approach consistently outperforms current state-of-the-art methods across different models in low query settings (nearly 70% fewer queries are used), a scenario especially relevant in real-world applications with limited queries and black-box access. Furthermore, by efficiently generating a suitable adversarial example, TEA provides an improved target initialization for established geometry-based attacks.
1 Introduction
@@
In black-box settings, targeted hard-label attacks have become an active area of research, with several techniques proposed in the literature, with state-of-the-art methods relying on the geometry of the decision boundary separating the source image from the target class. Geometry-informed attacks traverse in a lower-dimensional space and fall into two categories: Boundary Tracing Attacks and Gradient Estimation Attacks. Boundary Tracing Attacks ([3, 4, 18, 22]) perform walks along the decision boundary while Gradient Estimation Attacks ([6, 14, 16, 17, 19, 24, 25]) perform the same walks but using information about the approximate tangent/normal to the decision boundary in a local neighborhood to their adversarial image. Although powerful for local refinement, both approaches tend to burn through queries when the source image lies far from a given adversarial image in a target class, wasting many queries before reaching a narrow region where local geometry can more effectively be leveraged.
@@
Further, in practice, many real-world scenarios such as commercial pay-per-query APIs, impose severe constraints on the number of queries that can be made to a target model in a specified time. Often practical query limits may be on the order of a few hundred to fewer than a few thousand queries. Under these conditions, the limited feedback available makes it difficult to effectively use information about the local decision boundary geometry in the early stages of an attack. When the decision space is still wide, movement along restricted lower dimensional spaces leads to limited incremental progress, and gradient estimation methods often use queries in estimating local gradients by sampling predictions in a local neighborhood instead of moving. This raises a pressing need to develop methods that are efficient in the low-query regime. In a high query setting, this would also lead us to a good starting point to employ the geometry-informed methods since we arrive at a good adversarial point quickly and can use the geometry-informed information more effectively. To this end, we propose Targeted Edge-Informed Attack (TEA), a novel targeted adversarial attack designed for hard-label black-box scenarios within a restricted query budget. Rather than depending on the local properties of the decision boundary, TEA leverages the intrinsic features of the target image itself - specifically, the edge information obtained using Sobel filters [21] help identify prominent structural features in the target image. Edges encode high-magnitude spatial gradients that delineate object boundaries [21]. Research has shown that early layers fire on oriented edge filters [1, 26], and that shape/edge cues remain predictive even when textures are suppressed [10]. The core idea is to preserve these low-level features, while applying perturbations to the non-edge regions of an image, allowing us to stay in the target class while pushing the adversarial image towards the source image. When our progress plateaus, evidenced by a series of consecutive queries that fail to achieve further reduction in distance while maintaining target class prediction, one can switch to current state-of-the-art geometry-informed methods for a local refinement procedure. Hence we make the following contributions:
@@
Global initialization for low-query efficiency: We introduce an edge-informed perturbation strategy, enabling rapid progress toward the source image in the early stages of an adversarial attack. This strategy involves a two-step process: First, a global edge-informed search is performed, and then, edge-informed updates are applied to small patches using Gaussian weights.
@@
Empirical validation under strict query budgets: We perform extensive evaluations on the ImageNet validation dataset [7] across four architectures (ResNet-50 [12], ResNet-101 [12], VGG16 [20], and ViT [8]). Our attack consistently outperforms existing state-of-the-art hard-label methods, including HSJA, Adaptive History-driven Attack (AHA) [15], CGBA, and CGBA-H, under realistic query budgets (fewer than 1000 queries). To achieve a 60% reduction in distance from a target image to a source image, TEA required on average 251 queries across the four models - 70% fewer than AHA (the second fastest), which required 845 queries.
@@
The rest of this paper is structured as follows. Section 2 reviews related targeted hard-label adversarial attacks. Section 3 describes our methodology, while Section 4 presents experimental results. Finally, Section 5 discusses potential limitations and outlines directions for future research.
2 Related Works
@@
Incorporating historical query information, Adaptive History-driven Attack (AHA) [15] adapts its search trajectory based on previous successes and failures, while Decision-based query Efficient Adversarial Attack based on boundary Learning (DEAL) [6] employs an evolutionary strategy that concentrates queries on promising regions of the input space. SurFree [18] demonstrated that using 2D planes and semicircular trajectories toward the source image was an effective strategy. This was based on the fact that under the assumption of a flat decision boundary, the point on the boundary that is closest to the source image is precisely where the semicircular trajectory intersects it. Building on this idea, Curvature-Aware Geometric black-box Attack (CGBA) and its variant that is more suited for targeted attacks - CGBA-H [19], use normal estimation at the decision boundary to select the traversal plane. To the best of our knowledge, CGBA, and CGBA-H serve as the current state of the art in decision-based adversarial attacks.
3 Methodology
@@
Problem Statement. We consider a hard-label image classifier
@@
3.1 Global Edge-Informed Search
@@
[…]
@@
Global Interpolation. Starting from $x_{0}=x_{t}$, we perform iterative updates:
@@
where $\odot$ denotes the element-wise (Hadamard) product. The scaling factor $\alpha$ is optimized via a momentum based search. Masking out the edge regions in this update ensures that the interpolation primarily affects smooth areas, thus preserving the overall structure necessary for target classification. The complete procedure is summarized in Algorithm 2. Once the improvements stagnate, we transition to a local, patch-based refinement.
@@
3.2 Patch-Based Edge-Informed Search
After the global interpolation, some subregions of the image may still display significant discrepancies from $x_{s}$. In the patch-based refinement, we partition the image into patches
@@ 3.2 Patch-Based Edge-Informed Search
and for each patch, construct a local soft edge mask $M_{\mathcal{P}}$ analogously to $M_{\mathrm{edge}}$, but restricted to $\mathcal{P}$. Then, we apply a local interpolation:
@@
and search for the largest $\beta$ such that $\hat{y}(\widetilde{x}(\beta))=y_{t}$. Here, $G$ is a Gaussian weight applied across the patch to avoid artificial edges being created by our patches. This localized update is performed over multiple patches of varying sizes to steadily decrease the distance to the source image. The patch-based refinement is repeated until a termination criterion is met (e.g., 25 consecutive iterations with no further improvement). The corresponding illustration is depicted in Figure 1 and the relevant pseudocode is provided in Algorithm 3.
@@
Figure 2 offers a visual summary of TEA. The figure depicts the adversarial image throughout our method, and displays also a hotspot of individual pixel differences from the source image.
Figure 2: Visualization of TEA on a source–target image pair. The target image (initially classified as Bee-eater) is perturbed to resemble the source image (classified as Spoonbill), while preserving its original Bee-eater label. Global Edge-Informed Search efficiently applies edge-aware perturbations using only $15$ queries to achieve a $\approx$20reduction in distance to the source. Patch-Based Edge-Informed Search introduces localized, edge-aware modifications to small image regions, as seen in the hotspot of changes. Further refinement utilizing CGBA-H is illustrated in the narrow decision space.
@@
4 Experiments
In this section, we present our empirical evaluation on the ImageNet ILSVRC2012 dataset, benchmarking TEA against existing targeted hard-label attack strategies. In what follows, we describe our experimental setup and metrics, then detail the performance of each method under a range of query budgets. Our findings indicate that TEA achieves efficient distance reduction to the source image, particularly in the early query regime, before switching to the current state-of-the-art geometry-based attack CGBA-H for further refinement in a narrow decision space.
4.1 Setup and Metrics
@@
Dataset and Image Pairs. We sample 1000 source-target image pairs from the ImageNet ILSVRC2012 validation set, ensuring the two images in each pair come from distinct classes. All images are resized to $3\times 224\times 224$. Each pair $(x_{s},x_{t})$ contains a source image $x_{s}$, correctly classified under its label, and a target image $x_{t}$, also correctly classified under a different label. The attack goal is to modify $x_{t}$ to approach $x_{s}$ under the $\ell_{2}$ norm while keeping the prediction unchanged.
@@
Evaluation Metrics. Performance is quantified using three metrics. First, we compute the $\ell_{2}$ distance from $x_{s}$ to the adversarial example generated for each of the 1000 image pairs as queries progress. Second, the attack success rate (ASR) is defined as the fraction of image pairs for which an $\alpha\%$ reduction in $\|x_{s}-x_{t}\|$ is achieved within a given number of queries. Third, we also measure the performance of each method when there is a set fixed low-query budget by measuring the ASRs for all $\alpha$ when each method has used up $500$ queries. Finally, we integrate the $\ell^{2}$ distance versus queries curve to obtain the area under the curve (AUC), which provides an aggregated measure of how rapidly the distance is reduced.
Implementation Details. Our implementation employs a two-stage process. In the first stage, we perform our proposed methodology TEA, where edge-aware distortions are applied to the target image while preserving its classification, allowing a rapid reduction in distance to the source image. In the second stage, once the perturbation has brought the image into a narrow decision space, the method switches to CGBA-H for further refinement. The last query performed with TEA is referred to as the turning point throughout the study. We switch to CGBA-H since it consistently performs better than other methods in a high query setting across different architectures.
4.2 Results
Average $\ell_{2}$-Distance vs. Queries. Table 1 presents a comparative analysis of the median $\ell_{2}$ distances achieved in the low-query regime—specifically, within the range defined by the average turning point computed over one thousand image pairs. This analysis underscores the rapid decrease in perturbation norm facilitated by TEA during the early query stages. Figure 3 extends this evaluation by depicting the average percentage reduction in the $\ell_{2}$ distance against the queries used.
Table 1: Median $\ell_{2}$ distances computed until the turning point across different architectures. Lower $\ell^{2}$ values denote faster adversarial example generation.
@@ 4.2 Results
AUC Comparisons. We assess the efficiency of our aproach by computing the area under the median $\ell^{2}$-distance vs queries curve (AUC). Table 2 presents the AUC values for the low-query regime, up to the turning point. Across all architectures, TEA consistently yields lower AUC values.
Table 2: Average AUC values computed until the turning point across different architectures. Lower AUC values denote more effective early-stage distance reduction.
@@
High-Query Performance with CGBA-H refinement. Our hybrid TEA+CGBA-H strategy consistently matches or surpasses the state-of-the-art. Table 3 lists the median $\ell_{2}$ distances upto 20000 queries. Note that CGBA-H is randomness-dependent. Further, it doesn’t re-explore when trapped in a narrow decision space and its stability can be unpredictable and subject to local decision boundary geometry. Figure 7 illustrates the median percentage reduction in the distance between the images.
@@
Table 3: Median $\ell_{2}$ distances across different architectures. Here, TEA indicates using TEA until the turning point, and refining further with CGBA-H.

### v3

@@ Abstract
Deep neural networks for image classification remain vulnerable to adversarial examples — small, imperceptible perturbations that induce misclassifications. In black-box settings, where only the final prediction is accessible, crafting targeted attacks that aim to misclassify into a specific target class is particularly challenging due to narrow decision regions. Current state-of-the-art methods often exploit the geometric properties of the decision boundary separating a source image and a target image rather than incorporating information from the images themselves. In contrast, we propose Targeted Edge-informed Attack (TEA), a novel attack that utilizes edge information from the target image to carefully perturb it, thereby producing an adversarial image that is closer to the source image while still achieving the desired target classification. Our approach consistently outperforms current state-of-the-art methods across different models in low query settings (nearly 70% fewer queries are used), a scenario especially relevant in real-world applications with limited queries and black-box access. Furthermore, by efficiently generating a suitable adversarial example, TEA provides an improved target initialization for established geometry-based attacks.
I Introduction
@@
In black-box settings, targeted hard-label attacks have become an active area of research, with several techniques proposed in the literature [10, 11, 36, 8, 9], with state-of-the-art methods relying on the geometry of the decision boundary separating the source image from the target class. Geometry-informed attacks traverse in a lower-dimensional space and fall into two categories: Boundary Tracing Attacks and Gradient Estimation Attacks. Boundary Tracing Attacks ([3, 4, 28, 33]) perform walks along the decision boundary while Gradient Estimation Attacks ([7, 24, 26, 27, 30, 35, 37]) perform the same walks but using information about the approximate tangent/normal to the decision boundary in a local neighborhood to their adversarial image. Although powerful for local refinement, both approaches tend to burn through queries when the source image lies far from a given adversarial image in a target class, wasting many queries before reaching a narrow region where local geometry can more effectively be leveraged.
@@
Further, in practice, many real-world scenarios such as commercial pay-per-query APIs, impose severe constraints on the number of queries that can be made to a target model in a specified time. Often practical query limits may be on the order of a few hundred to fewer than a few thousand queries. Under these conditions, the limited feedback available makes it difficult to effectively use information about the local decision boundary geometry in the early stages of an attack. When the decision space is still wide, movement along restricted lower dimensional spaces leads to limited incremental progress, and gradient estimation methods often use queries in estimating local gradients by sampling predictions in a local neighborhood instead of moving. This raises a pressing need to develop methods that are efficient in the low-query regime. In a high query setting, this would also lead us to a good starting point to employ the geometry-informed methods since we arrive at a good adversarial point quickly and can use the geometry-informed information more effectively. To this end, we propose Targeted Edge-Informed Attack (TEA), a novel targeted adversarial attack designed for hard-label black-box scenarios within a restricted query budget. Rather than depending on the local properties of the decision boundary, TEA leverages the intrinsic features of the target image itself - specifically, the edge information obtained using Sobel filters [32] help identify prominent structural features in the target image. Edges encode high‑magnitude spatial gradients that delineate object boundaries [32]. Research has shown that early layers fire on oriented edge filters [1, 39], and that shape/edge cues remain predictive even when textures are suppressed [18]. The core idea is to preserve these low-level features, while applying perturbations to the non-edge regions of an image, allowing us to stay in the target class while pushing the adversarial image towards the source image. When our progress plateaus, evidenced by a series of consecutive queries that fail to achieve further reduction in distance while maintaining target class prediction, one can switch to current state-of-the-art geometry-informed methods for a local refinement procedure. Hence we make the following contributions:
@@
TEA: We introduce an edge-informed perturbation strategy, enabling rapid progress toward the source image in the early stages of an adversarial attack when limited queries are available. TEA involves a two-step process: First, a global edge-informed search is performed, and then, edge-informed updates are applied to small patches using Gaussian weights.
@@
Empirical validation under strict query budgets: We perform extensive evaluations on the ImageNet validation dataset [12] across four architectures (ResNet‑50 [20], ResNet‑101 [20], VGG16 [31], and ViT [13]) and an adversarially trained architecture (ResNet-50 [14]). Our attack consistently outperforms existing state‑of‑the‑art hard‑label methods, including HSJA, Adaptive History‑driven Attack (AHA) [25], CGBA, and CGBA‑H, under realistic query budgets (fewer than 1000 queries). To achieve a 60% reduction in distance from a target image to a source image, TEA required on average 251 queries across the four models - 70% fewer than AHA (the second fastest), which required 845 queries.
@@
The rest of this paper is structured as follows. Section 2 reviews related work on targeted hard-label adversarial attacks. Section 3 describes our methodology, while Section 4 presents experimental results. Finally, Section 5 outlines directions for future research.
II Related Works
@@
Incorporating historical query information, Adaptive History-driven Attack (AHA) [25] adapts its search trajectory based on previous successes and failures, while Decision-based query Efficient Adversarial Attack based on boundary Learning (DEAL) [7] employs an evolutionary strategy that concentrates queries on promising regions of the input space. SurFree [28] demonstrated that using 2D planes and semicircular trajectories toward the source image was an effective strategy. This was based on the fact that under the assumption of a flat decision boundary, the point on the boundary that is closest to the source image is precisely where the semicircular trajectory intersects it. Building on this idea, Curvature-Aware Geometric black-box Attack (CGBA) and its variant that is more suited for targeted attacks - CGBA-H [30], use normal estimation at the decision boundary to select the traversal plane. To the best of our knowledge, CGBA-H serves as the current state of the art in decision-based targeted adversarial attacks.
III Methodology
@@
Problem Statement. We consider a hard-label image classifier modeled by
@@
III-A Global Edge-Informed Search
@@
7:  $\text{edge\_mask}(i,j)\leftarrow\begin{cases}255,&\text{if }T_{\ell}\leq G(i,j)\leq T_{h},\\ 0,&\text{otherwise}\end{cases}$
8:  $\text{blurred}\leftarrow\text{GaussianBlur}(\text{edge\_mask},(b,b))$
9:  $\text{norm}\leftarrow\text{Normalize}(\text{blurred})$
@@
Global Interpolation. Starting from $x_{0}=x_{t}$, we perform iterative updates
@@
where $\odot$ denotes the element-wise (Hadamard) product. The scaling factor $\alpha$ is optimized via a momentum based search. Masking out the edge regions in this update ensures that the interpolation primarily affects smooth areas, thus preserving the overall structure necessary for target classification. The complete procedure is summarized in Algorithm 2. Once the improvements begin to stagnate, we transition to a local, patch-based refinement.
@@
III-B Patch-Based Edge-Informed Search
After the global interpolation, some subregions of the image may still display significant discrepancies from $x_{s}$. In the patch-based refinement, we then partition the image into randomly selected patches of random sizes
@@ III-B Patch-Based Edge-Informed Search
and for each patch, construct a local soft edge mask $M_{\mathcal{P}}$ by restricting $M_{\mathrm{edge}}$ to $\mathcal{P}$. Next, starting from our adversarial image $x_{k}$, we apply a local interpolation
@@
and search for the largest $\beta$ such that the classifier still predicts the target label, $\hat{y}(\widetilde{x}(\beta))=y_{t}$. Here, $G$ is a Gaussian weighting function over the patch that smoothly downweights updates near patch borders, helping to avoid artificial edges introduced by patch boundaries. We repeat this for different patches until a termination criterion is met (e.g., 25 consecutive iterations with no further improvement). The corresponding illustration is depicted in Figure 1 and the relevant pseudocode is provided in Algorithm 3.
@@
Figure 2 offers a visual overview of TEA. The figure depicts the adversarial image throughout our method, and displays a hotspot of individual pixel differences from the source image.
Figure 2: Visualization of TEA on a source–target image pair. The target image (initially classified as Bee-eater) is perturbed to resemble the source image (classified as Spoonbill), while preserving its original Bee-eater label. Global Edge-Informed Search efficiently applies edge-aware perturbations using only $15$ queries to achieve a $\approx$20% reduction in distance to the source. Patch-Based Edge-Informed Search introduces localized, edge-aware modifications to small image regions, as seen in the hotspot of changes. Further refinement utilizing CGBA-H is illustrated in the narrow decision space.
@@
IV Experiments
In this section, we present our empirical evaluation benchmarking TEA against existing targeted hard-label attacks. In what follows, we describe our experimental setup and metrics, then detail the performance of each method under a range of query budgets. Our findings indicate that TEA achieves efficient distance reduction to the source image, particularly in the early query regime, before switching to the current state-of-the-art geometry-based attack CGBA-H for further refinement in a narrow decision space.
IV-A Setup and Metrics
@@
Dataset and Image Pairs. We randomly sample 1000 source-target image pairs from the ImageNet ILSVRC2012 validation set, ensuring each pair contains images from distinct classes. All images are resized to $3\times 224\times 224$. Each pair $(x_{s},x_{t})$ contains a source image $x_{s}$, correctly classified under its label, and a target image $x_{t}$, also correctly classified under a different label. The goal is to modify $x_{t}$ to approach $x_{s}$ under the $\ell_{2}$ norm while keeping the prediction unchanged.
@@
Evaluation Metrics. Performance is quantified using three metrics. First, we compute the $\ell_{2}$ distance from $x_{s}$ to the adversarial example generated for each of the 1000 image pairs as queries progress. Second, the attack success rate (ASR) is defined as the fraction of image pairs for which an $\alpha\%$ reduction in $\ell_{2}$ distance between adversarial image and source image is achieved as compared to the target image and source image. Third, we also measure the performance of each method when there is a set fixed low-query budget by measuring the ASRs for all $\alpha$ when each method has used up $500$ queries. Finally, we integrate the $\ell_{2}$ distance versus queries curve to obtain the area under the curve (AUC), which provides an aggregated measure of how rapidly the distance is reduced.
Implementation Details. Our implementation employs a two-stage process. In the first stage, we perform our proposed methodology TEA, where edge-aware distortions are applied to the target image while preserving its target classification, allowing a rapid reduction in distance to the source image. In the second stage, once the perturbations have brought the image into a narrow decision space, the method switches to CGBA-H for further refinement (denoted as TEA throughout the study). The last query performed with TEA is referred to as the turning point throughout the study. We switch to CGBA-H since it consistently performs better than other methods in a high query setting across different architectures.
IV-B Results
TABLE I: Median $\ell_{2}$ distances computed until the turning point across different architectures.
Lower $\ell_{2}$ values denote faster adversarial example generation.
@@ IV-B Results
Average $\ell_{2}$-Distance vs. Queries. Table I presents a comparative analysis of the median $\ell_{2}$ distances achieved in the low-query regime—specifically, within the range defined by the average turning point computed over one thousand image pairs. This analysis underscores the rapid decrease in perturbation norm facilitated by TEA during the early query stages. Figure 3 extends this evaluation by depicting the average percentage reduction in $\ell_{2}$ distance against queries used. In the plots, the initial inflection point marks TEA’s transition from global exploration to patch-based refinement.
@@
AUC Comparisons. We assess the efficiency of our aproach by computing the area under the median $\ell_{2}$-distance vs queries curve (AUC). Table II presents the AUC values for the low-query regime, up to the turning point. Across all architectures, TEA consistently yields lower AUC values.
TABLE II: Average AUC values computed until the turning point across different architectures. Lower AUC values denote more effective early-stage distance reduction.
@@
Evaluation on an Adversarially Trained Model. In Figure 7, we present a comparative analyses of performance against an adversarial trained model (Resnet50 from MardyLab [14]) in the low-query region. We see that TEA consistently achieves more distortion, while also noting that all methods perform significantly worse compared to the standard Resnet50 model as seen in Figure 3.
Refer to caption Refer to caption
Figure 7: Comparison on an adversarially trained model. Left: median percentage decrease in $\ell_{2}$ distance against number of queries, with higher values indicating a more effective reduction method. Right: median $\ell_{2}$ distance against number of queries, with lower values indicating a more effective reduction method.
Edge Ablation Study To better understand the role of edge preservation in our method, we perform an ablation study by comparing three variants of TEA under identical settings. TEA perturbs non-edge pixels, preserving edge regions with a soft edge mask. INV-TEA inverts TEA’s soft edge mask to perturb primarily edge-like pixels. HALF-TEA perturbs non-edge and edge regions equally, using the same editable-pixel budget as TEA, and serves as a simple control baseline. Table III summarizes results in the low-query regime: median $\ell_{2}$ distance, ASR at $50\%$ distance reduction, and ASR at $75\%$ distance reduction. Across all four architectures, TEA achieves the lowest median $\ell_{2}$ and the highest success rates overall; HALF-TEA is consistently in between, and INV-TEA performs the worst, indicating that prioritizing non-edge regions proves crucial for early attack progress while preserving the semantic structure.
TABLE III: Edge ablation in the low-query regime.
Per architecture and budget we report median $\ell_{2}$ (lower is better), ASR at 50%, and ASR at 75% (higher is better).
Query ResNet-50 Query ResNet-101
$\ell_{2}$ ASR@50% ASR@75% $\ell_{2}$ ASR@50% ASR@75%
TEA Half Inv TEA Half Inv TEA Half Inv TEA Half Inv TEA Half Inv TEA Half Inv
100 75.1377 87.3465 87.5367 44.0 23.6 22.8 2.0 0.2 1.3 100 74.4637 86.1526 86.8610 44.4 24.1 23.2 2.8 0.3 1.0
200 61.8307 73.1697 73.3559 70.5 49.7 45.2 6.5 0.7 2.6 200 61.3517 71.7322 72.6071 72.4 53.3 47.3 6.9 0.9 3.1
300 55.3574 65.0474 66.5448 81.8 67.8 61.8 10.7 3.3 4.4 300 54.9709 63.5007 65.8087 83.9 69.7 62.0 11.8 3.2 4.7
400 51.8838 60.0257 63.2027 86.0 77.6 69.0 15.6 5.9 6.1 400 51.3084 58.5084 62.4517 87.5 77.6 69.0 17.1 6.7 6.0
500 49.8877 56.8181 61.6506 87.5 81.8 71.3 19.2 9.5 6.9 500 49.3153 55.3806 60.8018 89.6 82.6 72.2 19.7 9.5 6.5
Query VGG16 Query ViT
$\ell_{2}$ ASR@50% ASR@75% $\ell_{2}$ ASR@50% ASR@75%
TEA Half Inv TEA Half Inv TEA Half Inv TEA Half Inv TEA Half Inv TEA Half Inv
100 77.4206 90.7514 89.8802 38.1 17.6 20.8 2.0 0.3 0.8 100 67.4874 76.4322 77.6473 57.8 43.4 40.4 5.2 0.9 2.8
200 63.7679 75.3909 75.5661 66.3 45.0 40.8 5.8 0.7 2.3 200 56.2801 63.8529 66.0835 80.0 69.8 60.8 11.8 3.6 6.7
300 57.0986 66.5727 68.5222 78.7 64.8 56.9 10.8 2.9 3.9 300 50.8057 57.2091 60.8559 88.1 82.8 71.0 17.5 7.7 9.7
400 53.4352 61.0717 65.1434 83.3 74.9 62.9 16.8 7.4 5.0 400 47.9956 53.3073 58.3993 90.6 88.1 74.5 22.9 12.7 12.1
500 51.4759 57.5518 63.4061 84.7 79.2 66.6 19.5 11.4 6.8 500 46.3913 51.0218 57.2895 92.1 90.6 76.0 26.1 16.2 13.0
Randomness and Stability Across Patch Selection Our patch-based refinement uses random patch locations and sizes, introducing stochasticity. To quantify this effect, we executed five independent runs per model. For each run and query budget, we computed the average $\ell_{2}$ distance across all pairs; we then summarized the five runs by reporting mean $\pm$ standard deviation (std) of the average $\ell_{2}$ distance across all pairs. As seen in Table IV, across all models and budgets, variability is small relative to the mean. The standard deviation is typically below $0.3$ (well under $1\%$ of the mean).
TABLE IV: Randomness analysis across five runs: $\ell_{2}$ (mean $\pm$ std).
Budget ResNet50 ResNet101 VGG16 ViT
50 88.1415 $\pm$ 0.0254 87.0992 $\pm$ 0.0637 90.6675 $\pm$ 0.0553 78.7152 $\pm$ 0.0277
100 75.2155 $\pm$ 0.2338 74.5036 $\pm$ 0.0676 77.2608 $\pm$ 0.1085 67.5654 $\pm$ 0.0652
150 67.3196 $\pm$ 0.2301 66.6368 $\pm$ 0.1022 69.1007 $\pm$ 0.1479 60.6342 $\pm$ 0.1187
200 62.0338 $\pm$ 0.2607 61.3794 $\pm$ 0.1318 63.6530 $\pm$ 0.1896 56.1255 $\pm$ 0.0829
250 58.2941 $\pm$ 0.2458 57.6515 $\pm$ 0.1356 59.7896 $\pm$ 0.2565 53.0414 $\pm$ 0.0537
300 55.5589 $\pm$ 0.1995 54.9321 $\pm$ 0.1397 56.9954 $\pm$ 0.2654 50.8059 $\pm$ 0.0627
400 51.9530 $\pm$ 0.1907 51.3780 $\pm$ 0.1564 53.4018 $\pm$ 0.2513 47.8666 $\pm$ 0.0654
500 49.8352 $\pm$ 0.1522 49.3863 $\pm$ 0.1306 51.3763 $\pm$ 0.2361 46.2528 $\pm$ 0.0899
Lower Resolution Performance We evaluate TEA, CGBA, CGBA-H and AHA at two lower input sizes, $128\times 128$ and $64\times 64$. The Tables V and VI report the median $\ell_{2}$ distance to the source image achieved at fixed query budgets; lower is better. VGG-16 and ViT required standard architectural adjustments at reduced dimensions.
TABLE V: Median $\ell_{2}$ distances at $\times128$.
Query ResNet50 Query ResNet101
AHA CGBA CGBA-H TEA AHA CGBA CGBA-H TEA
50 52.812610 56.892204 56.892204 53.891792 50 52.325765 55.829930 55.830184 52.836270
100 49.792591 52.638543 52.610928 44.656296 100 49.256166 51.690833 52.297024 43.908848
150 47.580821 48.751302 48.900604 40.200085 150 47.165379 47.665892 48.747030 39.356266
200 45.816245 47.670539 47.708520 37.588080 200 45.418667 46.499661 48.731472 36.744068
250 44.242092 47.291906 47.416906 35.637354 250 43.904684 45.875023 48.286582 34.803333
300 42.705044 45.310037 45.394565 34.083014 300 43.015997 43.947135 46.251110 33.364790
350 41.704023 45.917060 46.053402 32.915845 350 42.285534 45.637712 48.112825 32.363468
400 40.606535 43.876574 44.083020 31.711626 400 41.279571 43.208419 47.001603 31.200894
Query VGG16 Query ViT
AHA CGBA CGBA-H TEA AHA CGBA CGBA-H TEA
50 53.742676 57.595045 57.595044 54.653168 50 49.717520 61.999604 61.999605 51.518260
100 50.687319 53.132007 53.124912 45.101308 100 46.560914 56.845693 56.760291 42.468795
150 48.452131 48.769476 48.688020 40.641963 150 44.133746 51.938031 51.849316 37.765469
200 46.376375 47.648656 47.548296 37.729822 200 42.135314 52.195369 52.190025 34.882641
250 44.719990 46.665404 46.692002 35.775914 250 40.692779 50.410471 50.458582 32.847132
300 43.824315 44.433300 44.557708 34.574623 300 39.469095 47.456320 47.224463 31.225525
350 42.589694 45.370771 45.405220 33.251792 350 38.537397 49.233241 49.101166 30.127403
400 41.597263 43.446274 43.461510 32.151386 400 37.460420 47.532119 47.482755 29.117229
TABLE VI: Median $\ell_{2}$ distances at $\times64$.
Query ResNet50 Query ResNet101
AHA CGBA CGBA-H TEA AHA CGBA CGBA-H TEA
50 28.562640 31.263057 31.263184 29.709433 50 28.527201 31.240249 31.240327 29.828940
100 26.583674 28.666067 28.734703 25.884418 100 26.545651 28.491136 28.543173 26.137666
150 25.094041 26.041276 26.107465 23.768843 150 25.101155 25.824065 25.971276 23.934921
200 23.449901 24.898503 25.231694 21.971898 200 23.850520 24.798062 25.058821 22.132178
250 22.191710 24.117243 24.478960 20.671545 250 22.805084 23.948980 24.090509 20.863024
Query VGG16 Query ViT
AHA CGBA CGBA-H TEA AHA CGBA CGBA-H TEA
50 28.318550 30.844857 30.837613 29.530338 50 30.081730 32.758422 32.758422 31.046581
100 26.390916 28.302722 28.346767 25.606639 100 27.818672 29.806229 29.807765 27.077744
150 24.731502 25.578313 25.651390 23.405260 150 26.072833 26.316862 26.340219 24.609781
200 23.409042 24.719593 24.810609 21.847981 200 24.510567 24.761248 24.786895 22.262164
250 22.453166 23.539063 23.687779 20.527478 250 23.248782 23.717796 23.829428 20.461638
As observable in the results, at lower input resolution sizes, each patch modification distorts a larger fraction of the target image and thus contributes more to the overall $\ell_{2}$-norm reduction. Consequently, maintaining the adversariality of the image becomes much harder, and TEA therefore terminates earlier than for higher resolutions.
Perceptual Metrics While adversarial examples are generated to trick ML models, it is worth considering their impact on human perception [5, 17]. We report SSIM [38] and FSIM [40] alongside distortion at a fixed budget of 400 queries. As shown in Table VII, TEA achieves high distortion while slightly improving FSIM. A small drop in SSIM is expected, since TEA’s perturbations are more structured. Prior image quality assessment (IQA) studies [40, 16] indicate FSIM correlates more strongly with human perception than SSIM, which helps contextualise these differences.
TABLE VII: SSIM and FSIM at 400 queries on 1000 source—target pairs.
Lower $\ell_{2}$ indicates smaller perturbations; higher SSIM/FSIM indicates greater perceptual similarity to the source image.
ResNet50 ResNet101
Metric CGBA CGBA-H AHA TEA CGBA CGBA-H AHA TEA
$\ell_{2}$ 85.579 71.479 64.172 51.9530 82.483 68.803 67.051 51.3780
SSIM 0.531765 0.530459 0.231784 0.478806 0.525847 0.528782 0.223511 0.491950
FSIM 0.251337 0.252585 0.232117 0.294672 0.252746 0.251235 0.223831 0.288153
VGG16 ViT
Metric CGBA CGBA-H AHA TEA CGBA CGBA-H AHA TEA
$\ell_{2}$ 89.880 70.559 68.538 53.4018 72.839 61.300 59.857 47.8666
SSIM 0.534783 0.532189 0.231784 0.457879 0.580527 0.582544 0.201767 0.548110
FSIM 0.251129 0.251239 0.232117 0.305919 0.220927 0.220213 0.202171 0.257571
@@
High‐Query Performance with CGBA-H refinement. Our hybrid TEA+CGBA‑H strategy consistently matches or surpasses the state-of-the-art. Table VIII lists the median $\ell_{2}$ distances upto 20000 queries. Note that since CGBA-H is randomness-dependent, it doesn’t re-explore when trapped in a narrow decision space and its stability is subject to local decision boundary geometry. Figure 8 illustrates the median percentage reduction in the distance between the images.
TABLE VIII: Median $\ell_{2}$ distances across different architectures.
@@
Here, TEA indicates using TEA until the turning point, and refining further with CGBA-H.